OpenStudy (anonymous):

If f(x) = x2 and g(x) = x + 6, find g(f(0)) please help me :)

OpenStudy (anonymous):

anyone?????

OpenStudy (freckles):

f(0) means replace x with 0 in f(x)=x^2

OpenStudy (freckles):

once you do that we will move on the plug in whatever results from there into g(x)=x+6

OpenStudy (anonymous):

\(g(f(x))\) at \(x=0\) is what is being asked so plug in \(0\) for \(x\):\[g(f(0))\]what is \(f(0)\):\[f(0)=(0)^2\]\[f(0)=0\]replace this value in \(g(f(0))\):\[g(f(0))=g(0)\]now solve \(g(0)\):\[g(0)=(0)+6\]\[g(0)=6\] \[\text{OR}\] the question asks to find the function of \(g\) of \(f\) of \(x\) so everywhere in \(g(x)\) that you see an \(x\) you would replace it with \(f(x)\):\[g(x)=x+6\]\[g(f(x))=f(x)+6\]\(f(x)=x^2\) so:\[g(f(x))=x^2+6\]now plug in \(0\) as the value of \(x\) to find \(g(f(x))\):\[g(f(0))=(0)^2+6\]\[g(f(0))=0+6\]\[{g(f(0))=6}\]